# Unleashing the Power of Meta-tuning for Few-shot Generalization Through Sparse Interpolated Experts

### v1

@@ 5 Experiment
Experimental setup.  We conduct meta-tuning experiments on Meta-dataset (MD(Triantafillou et al., 2020), which is a widely studied large-scale cross-domain few-shot learning benchmark. As in PMF (Hu et al., 2022), we adhere to the official guidelines and employ the standard meta-training and meta-testing splits for meta-tuning and meta-testing. We select all hyperparameters and the meta-tuned checkpoint for testing using the official meta-validation split. In order to obtain a more comprehensive evaluation of the meta-tuned models, we introduce additional OOD datasets for meta-testing only, which were not used during the meta-tuning process on MD.
@@
Evaluation.  At meta-testing time, we resort to the ProtoNet (Snell et al., 2017) classifier for performing direct inference on each few-shot testing tasks without further adaptation. When considering task-specific fine-tuning on the support sets, we follow the same protocols in PMF (Hu et al., 2022) for all models, using Full fine-tuning (fine-tuning the entire model) and LoRA (Hu et al., 2021) as fine-tuning methods. Namely, for each dataset, we perform a hyperparameter search on a small set of validation tasks to obtain the optimal learning rate for task-specific fine-tuning using the Adam optimizer for 50 steps for each testing tasks from that dataset.
Table 1: Few-shot testing results on the Meta-dataset benchmark and additional OOD testing datasets for methods using DINO-ViT-Small backbone. ${}^{\dagger}$ and ${}^{\ddagger}$ respectively indicate published results in ${}^{\dagger}$(Hu et al., 2022) and ${}^{\ddagger}$(Basu et al., 2023). Gray indicates our method.
w/o fine-tuning with gradient-based fine-tuning
Datasets ${}^{\dagger}$re ${}^{\dagger}$ SoftMerge SMAT ${}^{\ddagger}$re+full ${}^{\dagger}$+full SoftMerge+full SMAT+full ${}^{\ddagger}$re+LoRA PMF+LoRA SMAT+LoRA
ImageNet 73.48 73.54 74.33 74.94 73.54 74.59 74.71 75.24 74.22 73.54 75.72
Aircraft 62.17 88.33 88.80 89.49 75.4 88.33 90.6 90.78 80.8 89.75 90.71
Omniglot 54.33 91.79 91.24 89.54 78.7 91.79 92.01 90.83 80.8 92.78 90.99
CUB 85.37 91.02 91.54 92.48 85.4 91.02 91.95 92.48 85.8 91.17 92.42
DTD 83.67 81.64 80.98 85.86 86.9 86.61 86.84 88.34 86.8 86.73 88.28
Quickdraw 60.59 79.23 78.98 78.83 73.6 79.23 79.90 78.83 72.7 79.23 78.83
Fungi 56.26 74.2 72.40 72.8 54.7 74.20 72.40 72.80 59.8 75.44 72.80
VGGFlower 94.45 94.12 96.89 97.19 94.2 94.12 97.01 97.19 94.8 96.05 97.25
ID Avg 71.29 84.23 84.40 85.14 77.81 84.99 85.56 85.81 79.47 85.59 85.88
TrafficSig 53.7 54.37 56.21 58.51 87.3 88.85 89.91 90.83 88.1 89.14 90.18
MSCOCO 54.58 57.04 55.75 57.35 61.5 62.59 62.15 63.07 62.1 61.71 63.38
Cifar10 85.64 80.82 84.58 83.95 92.48 89.61 91.84 92.08 93.33 91.53 92.46
Cifar100 76.86 69.11 70.85 74.85 86.13 82.54 85.88 85.91 86.17 85.06 85.88
MNIST 78.57 93.33 94.16 94.53 92.54 96.44 96.20 96.73 94.98 96.41 96.46
Sketch 47.25 41.10 43.30 48.91 56.39 49.65 53.85 56.55 57.34 47.59 55.63
Food 91.73 91.37 89.84 92.31 92.03 91.73 90.48 92.31 92.06 92.01 92.31
Clipart 55.19 53.92 54.83 59.87 67.18 62.83 65.50 65.76 66.51 60.6 66.07
Pet 62.64 61.89 63.04 65.59 65.08 62.97 63.36 67.43 65.06 62.71 67.77
Cars 34.58 38.00 36.21 36.79 40.98 40.07 41.62 42.39 39.49 42.37 40.05
OOD Avg 64.07 64.10 64.87 67.27 74.16 72.73 74.08 75.31 74.51 72.91 75.02
5.1 SMAT achieves new SOTA performance.

### v3

@@ 2 Related Work
3 Preliminary: Meta-tuning
Meta-tuning aims to improve the few-show learning performance of a pre-trained model on downstream few-shot testing tasks - usually by directly meta-training the pre-trained model over a collection of labeled few-shot training task episodes  (Hu et al., 2022; Min et al., 2021). Specifically, we assume the availability of a pre-trained model $f_{\bm{\theta}^{\mathtt{pre}}}$ as initialization for meta-training, and a training task distribution ${\mathcal{P}}_{ID}({\mathcal{T}})$ from which we may sample fully labeled few-shot training tasks ${\mathcal{T}}_{i}\sim{\mathcal{P}}_{ID}({\mathcal{T}})$. Note that we explicitly denote this as in-distribution ($ID$). In particular, in the supervised setting, each training task ${\mathcal{T}}_{i}$ takes the form of ${\mathcal{T}}_{i}:=\{{\mathcal{L}}_{i},{\mathcal{T}}_{i}^{s},{\mathcal{T}}_{i}^{q}\}$, where ${\mathcal{L}}_{i}$ is the task loss to be minimized, ${\mathcal{T}}_{i}^{s}:=\{\mathbf{x}_{i,j}^{s},\mathbf{y}_{i,j}^{s}\}_{j=1}^{N_{i}^{s}}$ and ${\mathcal{T}}_{i}^{q}\!:=\!\{\mathbf{x}_{i,j}^{q},\mathbf{y}_{i,j}^{q}\}_{j=1}^{N_{i}^{q}}$ are labeled support and query sets of $N_{i}^{s}$, $N_{i}^{q}$ input-target pairs, respectively. We use the shorthand notations $\mathbf{X}$ and $\mathbf{Y}$ to represent a set of inputs and labels, respectively. Meta-tuning is then realized through the typical episodic-learning setting familiar from meta-learning, i.e., the direct optimization of $\bm{\theta}$ on the few-shot learning objective which considers minimizing the task loss on the query predictions given information of the support set, i.e., $\bm{\theta}^{*}:=\operatorname*{arg\,min}_{\bm{\theta}\in\Theta}\mathbb{E}_{{\mathcal{P}}_{ID}}[{\mathcal{L}}_{i}(f_{\bm{\theta}}(\mathbf{X}^{q}_{i},{\mathcal{T}}^{s}_{i}),\mathbf{Y}^{q}_{i})]$, over training task episodes sampled from ${\mathcal{P}}_{ID}({\mathcal{T}})$. At test time, we expect to encounter both ID and OOD testing tasks i.e., $\tilde{{\mathcal{T}}}:=\{\tilde{{\mathcal{T}}}^{s}_{i},\tilde{\mathbf{X}}_{i}^{q}\}\sim{\mathcal{P}}_{ID}\cup{\mathcal{P}}_{OOD}$, where ${\mathcal{P}}_{OOD}\neq{\mathcal{P}}_{ID}$ is an unseen OOD task distribution. For each testing task, we evaluate the few-shot generalization performance of the meta-tuned model by predicting query labels. Hence, our objective is to develop a meta-tuning algorithm that enables the meta-tuned $\bm{\theta}^{*}$ to attain optimal few-shot generalization performance across both ID and OOD testing tasks.
Figure 2: Overview of the proposed Sparse MetA-Tuning approach, showing meta-training and inference procedures for a single task ${\mathcal{T}}_{i}$. SMAT meta-learns a shared knowledge pool $\mathcal{M}$ consisting of $|\mathcal{M}|$ sparse interpolated experts characterized by a common, learnable set of dense parameters $\bm{\theta}^{\delta}$ and distinct, learnable sets of gating masks $\{\bm{z}_{m}\}^{|\mathcal{M}|}_{m=1}$ with sparsity constraints. To construct each task-specific model $\bm{\theta}_{i}$ for both meta-training and inference, (1) SMAT first combines experts via a weighted-sum with merging weights $\bm{\alpha}_{i}$ generated by a meta-learned hypernetwork $h_{\bm{\zeta}}$ based on the task’s support set $\mathcal{T}_{i}^{s}$. (2) The experts are then subsequently combined with the frozen pre-trained model $\bm{\theta}^{\mathtt{pre}}$ to enhance both in-distribution (ID) and out-of-distribution (OOD) generalization performance. Alongside (3) the query prediction loss ${\mathcal{L}}^{\mathtt{ce}}_{\mathcal{T}_{i}^{q}}(\bm{\theta}_{i})$, (4) knowledge distillation with task-specific dense teachers ${\mathcal{L}}^{\mathtt{kd}}_{\mathcal{T}_{i}^{q}}(\bm{\theta}_{i},\bm{\theta}_{i}^{\mathtt{tr}})$ is introduced during meta-training to promote specialization and cooperation of the sparse interpolated experts, ensuring optimization success.
4 SMAT: Sparse MetA-Tuning
4.1 Meta-training
As discussed in Section 1, naively sharing and updating all pre-trained parameters across all tasks in meta-tuning leads to task interference in optimization (Yu et al., 2020; Wang et al., 2020). To address this issue, we instead hypothesize that the solution for each task (ID or OOD) comprises a task-specific mixture of a common pool of knowledge covering a broad range of tasks. The knowledge pool is represented by distinct sets of model parameters (i.e., experts), which can be combined cooperatively as a complementary addition to the pre-trained model to promote systematic generalization. Formally, we assume that each task-specific model $\bm{\theta}_{i}$ is derived from aggregating the experts via a task-specific weighted sum in the parameter space:
where ${\bm{\theta}}_{m}^{\delta}$ represents the $m$-th expert in the pool ${\mathcal{M}}$, and $\bm{\alpha}_{i}:=[\alpha_{i,1},\alpha_{i,2},...,\alpha_{i,|{\mathcal{M}}|}]$ are the weights. This way of merging experts in Eqn. (1) is appealing due to its high expressiveness, as supported by recent findings revealing that merging multiple sets of specialized parameters through simple arithmetic can lead to a better overall multi-task solution (Ilharco et al., 2023; Matena & Raffel, 2021). Moreover, the complementary addition of experts to the pre-trained backbone effectively decouples the two for optimization - meta-training $\{\bm{\theta}^{\delta}_{m}\}^{|{\mathcal{M}}|}$ while keeping $\bm{\theta}^{\mathtt{pre}}$ frozen provides sufficient capacity and preserves pre-trained knowledge.
That said, meta-tuning, in our case, is seen as a procedure for discovering a generalizable selection rule for assigning appropriate experts to tasks. We present an overview of our proposed method in Fig. 2.
Sparsification of experts.    At the core of our method lies the question on where and how to find a set of experts in the model during meta-tuning. Previous studies on partitioned meta-learning, as well as recent works on Mixture-of-Experts, propose a seemingly reasonable solution: Experts are incorporated into a heuristically (and a-priori) chosen subset of modules in the model. For example, the batch-norm (BN) layers in ResNets (Triantafillou et al., 2021b) or the MLP layers in Vision Transformers (ViT) (Puigcerver et al., 2023). Although these approaches based on fixed partitioning are more memory-efficient and generally result in improved performance compared to their non-expert counterparts, they suffer from a significant bias due to their manually crafted selection (i.e., the concrete choice of $\bm{\theta}_{m}^{\delta}$). This bias may be suboptimal when it conflicts with the intricate dynamics of meta-training and, in addition, is not always model-agnostic, e.g., there are no BN layers in ViT.
Thus, instead of specifying experts as prior knowledge, we propose to automatically identify the sparsity patterns in experts through meta-tuning using a maximum likelihood objective while considering sparsity constraints on the experts’ capacity, thus encouraging specialization. Meta-learning sparsity patterns not only shows minimal bias but also allows for differentiation in the architecture of experts, giving our model a wider range of capabilities to better handle various types of distribution shifts. (Lee et al., 2022).
Specifically, we induce a learnable sparsity pattern in the $m$-th expert in the form of a sparse reparameterization $\bm{\theta}^{\delta}_{m}\odot\bm{z}_{m}$ of the dense expert, where $\bm{z}_{m}$ denotes a learnable binary mask with dimension $|\bm{z}_{m}|=|\bm{\theta}^{\delta}_{m}|$, and $\odot$ denotes element-wise multiplication. To learn binary masks through gradient-based optimization, we take inspiration from (Schwarz & Teh, 2022), and reparameterize the $m$-th binary mask using an underlying continuous distribution $q_{\bm{\phi}_{m}}$ with parameters $\bm{\phi}_{m}$. The reparameterization samples, i.e., $\mathbf{s}_{m}\!\sim\!q_{\bm{\phi}_{m}}(\mathbf{s})$, can be transformed to have values exactly 00 and $1$ through a hard rectification $\mathbf{z}_{m}\!=\!g(\mathbf{s}_{m})\!:=\!\min(1,\max(0,\mathbf{s}_{m}))$. As a result, sparsity in $\bm{z}_{m}$ can be enforced by limiting the probability of $\bm{s}_{m}$ being non-zero which can be straight-forwardly expressed through the CDF $Q$ of $q_{\bm{\phi}_{m}}$, i.e. $1-Q_{\bm{\phi}_{m}}(\mathbf{s}\leq 0)$. Choosing $q_{\bm{\phi}_{m}}$ as the stretched hard concrete distribution (Louizos et al., 2018) enables both gradient-based optimization through reparameterization as well as analytical evaluation of the CDF.
From sparse experts to sparse interpolated experts.    Though the aforementioned sparse reparameterization eventually reduces parameter counts (e.g., by removing zeros in expert parameters), it leads to a strong increase in the number of parameters compared to its non-sparse counterpart at the beginning of meta-tuning – due to learning $\bm{z}_{m}$ and $\bm{\theta}_{m}^{\delta}$ concurrently. Moreover, assigning each expert its own dense underlying parameters $\bm{\theta}_{m}^{\delta}$ greatly hinders knowledge transfer among experts, which contradicts the principles of partitioned meta-learning emphasizing collaboration of both task-agnostic and task-specific components.
To tackle these issues, we propose sharing dense parameters across sparse reparameterization of different experts, i.e., $\bm{\theta}_{m}^{\delta}=\bm{\theta}^{\delta},\forall m$. By rearranging Eqn. (1) slightly, the experts now essentially become different sparse interpolations between the same pre-trained and meta-tuned models (more details in Appendix D). Although this formulation of MoE may initially appear bold, it is well-supported by recent works (Panigrahi et al., 2023b) which suggest that multiple task-specific optimal points can coexist between the same set of pre-trained and meta-tuned models, offering favorable trade-offs for both in-distribution (ID) and some out-of-distribution (OOD) performance. Therefore, our approach of learning sparse interpolated experts can be seen as an inductive bias that promotes the recovery of these optimal interpolation points through meta-tuning. Favorably, this partitioning allows for knowledge transfer among experts through the overlapping regions in their masks.
Meta-learning expert selection through a hypernetwork.     In theory, there are two possible approaches to achieve task-specific inference for the expert merging scores $\bm{\alpha}_{i}$. The first one is a meta-learned merging score initialization $\bm{\alpha}_{0}$ combined with inner-loop gradient-descent on the task support set; the second involves a meta-learned global hypernetwork, $h_{\bm{\zeta}}({\mathcal{T}}_{i}^{s})$, parameterized by $\bm{\zeta}$, that directly outputs $\bm{\alpha}_{i}$ conditioned on the task support set ${\mathcal{T}}_{i}^{s}$. We opt for the latter approach as it scales better with larger model sizes. For implementation, we use the pre-trained model $f_{\bm{\theta}^{\mathtt{pre}}}$ to encode each support image $\mathbf{x}_{i,j}^{s}\in\mathcal{T}_{i}^{s}$ into a vector embedding $f_{\bm{\theta}^{\mathtt{pre}}}(\mathbf{x}_{i,j}^{s})$. The support set embeddings are aggregated into class prototypes which are then concatenated into a sequence and fed into a single trainable transformer block to obtain $\bm{\alpha}_{i}^{\prime}\in\mathbb{R}^{|{\mathcal{M}}|}$ as the output. We treat $\bm{\alpha}_{i}^{\prime}$ as the logits for activating the experts. Instead of choosing the top-k activated ones, which can cause training instability issues and, more importantly, restrict the number of experts per task, we employ the Gumbel-Sigmoid trick (Jang et al., 2017) to sample a soft activation value $\in(0,1)$ for each expert, followed by normalization to obtain $\bm{\alpha}_{i}\in(0,1)^{|{\mathcal{M}}|}$.
Enhancing expert specialization through task-specific dense teachers.    Specialization and cooperation among experts play a crucial role in a MoE model. One way to promote specialization is by penalizing the similarity between experts. For example, an orthogonal penalty can be applied to pairs of experts. However, incorporating such explicit penalties makes the optimization problem in the meta-objective more challenging, as it introduces a trade-off with respect to the few-shot prediction performance. To circumvent this trade-off, we propose an alternative approach that utilizes a knowledge distillation loss $\mathcal{L}^{\mathtt{kd}}_{\mathcal{T}_{i}^{q}}(\bm{\theta}_{i},\bm{\theta}_{i}^{\mathtt{tr}})$ (Hinton et al., 2015) between the merged MoE model $\bm{\theta}_{i}$ and a teacher network $\bm{\theta}_{i}^{\mathtt{tr}}$. By using a highly task-specific teacher, the distillation loss imposes the weighted-sum of experts that constitute each $\bm{\theta}_{i}$ to mimic the behaviour of the teacher, thus enhancing both specialization and cooperation implicitly through knowledge transfer. Furthermore, the predictive performance of $\bm{\theta}_{i}$ is not impeded as the loss consistently encourages the current student to achieve a better task-specific generalization on the query set, similar to that of the teacher.
We generate the teacher $\bm{\theta}_{i}^{\mathtt{tr}}$ dynamically in each meta-training episode by performing $K$-step gradient descent starting from $\bm{\theta}_{i}$ on the query loss $\mathcal{L}^{\mathtt{ce}}_{\mathcal{T}_{i}^{q}}$ (we find $K=1$ is sufficient in practice). It is worth noting that unlike $\bm{\theta}_{i}$, we do not limit the capacity of $\bm{\theta}_{i}^{\mathtt{tr}}$ through sparse regularization. This results in a teacher model that is both dense in modulation (i.e., $\bm{\theta}_{i}^{\mathtt{tr}}-\bm{\theta}^{\mathtt{pre}}$), making it expressive and highly task-specific. Importantly, we do not propagate the loss gradients through $\bm{\theta}_{i}^{\mathtt{tr}}\rightarrow\bm{\theta}_{i}$. This approach is thus similar to the use of bootstrapping in optimization-based meta-learning (Flennerhag et al., 2022; Tack et al., 2024).
$\min\limits_{\bm{\theta}^{\delta},\bm{\zeta},\bm{\Phi}}\max\limits_{\bm{\lambda}\geq\mathbf{0}}$ $\mathbb{E}_{{\mathcal{T}}_{i}\sim{\mathcal{P}}_{ID}}\big{[}{\mathcal{L}}^{\mathtt{ce}}_{\mathcal{T}_{i}^{q}}(\bm{\theta}_{i})+\mathcal{L}^{\mathtt{kd}}_{\mathcal{T}_{i}^{q}}(\bm{\theta}_{i},\bm{\theta}_{i}^{\mathtt{tr}})\big{]}$
$+\sum_{m=1}^{|{\mathcal{M}}|}\lambda_{m}(\frac{1}{|\bm{\phi}_{m}|}\sum_{k=1}^{|\bm{\phi}_{m}|}\tau-Q_{\bm{\phi}_{m}}(s_{k}\leq 0)),$
where $\bm{\theta}_{i}=\bm{\theta}^{\mathtt{pre}}+\sum_{m=1}^{|{\mathcal{M}}|}\alpha_{i,m}(\bm{z}_{m}\odot\bm{\theta}^{\delta}),$
$\bm{z}_{m}\sim q_{\bm{\phi}_{m}};\bm{\alpha}_{i}\sim h_{\bm{\zeta}}({\mathcal{T}}_{i}^{s}),$ (2)
in which the objective is the minimization problem in the first line while the aforementioned sparsity constraints translate to the maximization problem in the second. We put the sparsity constraints on individual masks, hence experts, as opposed to the overall $\bm{\theta}_{i}$ to allow task-dependent model capacity in $\bm{\theta}_{i}$ through selective merging. For simplicity, we set a common constraint for all masks i.e., $\tau_{m}=\tau,\forall m$, which is treated as a hyperparameter. We use simultaneous gradient descent and projected gradient ascent for optimizing Eqn. (4.1). To avoid over-penalizing the model capacity from surpassing the sparsity constraints, we reset $\lambda_{m}$ to zero after its associated sparsity constraint is met (Gallego-Posada et al., 2022). This results in final sparsity close to the target $\tau$, enabling precise control of the expert sparsity levels. The pseudocodes for meta-training can be found in Section A.1.
4.2 Meta-testing
In the next two sections, we outline the procedures for fine-tuning meta-tuned models using SMAT. SMAT is fully compatible with existing off-the-shelf fine-tuning techniques. Additionally, we introduce a gradient-free fine-tuning method specifically designed for SMAT, which can be optionally employed during downstream task adaptation at test time in more computation-restricted scenarios. Our pseudocodes and ablation results for various fine-tuning techniques with SMAT are provided in Section A.2 and Section C.2 .
Gradient-free optimization for expert selection.    Prior work has discovered that further task-specific adaptation of a meta-trained model is essential for improving its performance on OOD tasks during meta-testing (Hu et al., 2022; Chen et al., 2023; Li et al., 2022). Although effective, one major limitation of these adaptation methods is the reliance on back-propagation of the gradients, which can be expensive, making these methods inefficient and potentially impractical due to poor scaling with model size. To this end, we propose an adaptation strategy, specifically designed for SMAT, that bypasses the gradient computation. At the core of our method lies the hypothesis that each expert selection score can be descreatzed i.e, either 0 or 1, which aligns with the intuition that each expert knowledge is either required or not for solving each tasks. We can thus optimize the expert selection score in a binary hypothesize space i.e, $\bm{\alpha}_{i}\in\{0,1\}^{|{\mathcal{M}}|}$.
Gradient-based fine-tuning with SMAT.    Following (Hu et al., 2022), we also consider gradient-based fine-tuning of our meta-tuned model at meta-testing time. With only few changes, our method is fully compatible with existing full fine-tuning (i.e., fine-tuning the entire model), and parameter-efficient fine-tuning (PEFT) techniques. Specifically, we use $\bm{\theta}_{i}$ in Eqn. (4.1) as the task-specific model initialization, before applying any off-the-shelf fine-tuning technique for further optimizing $\bm{\theta}_{i}$ on the support set of each task.
@@ 5 Experiment
Setup.  We conduct meta-tuning experiments on Meta-dataset (MD(Triantafillou et al., 2020), which is a widely studied large-scale cross-domain few-shot learning benchmark. As in PMF (Hu et al., 2022), we adhere to the official guidelines and employ the standard meta-training and meta-testing splits for meta-tuning and meta-testing. We select all hyperparameters and the meta-tuned checkpoint for testing using the official meta-validation split. In order to obtain a more comprehensive evaluation of the meta-tuned models, we introduce additional OOD datasets for meta-testing only, which were not used during the meta-tuning process on MD.
Table 1: Few-shot testing results on the Meta-dataset benchmark and additional OOD testing datasets for methods using DINO-ViT-Small backbone. and respectively indicate published results in (Hu et al., 2022) and (Basu et al., 2023). Gray indicates our method.
w/o fine-tuning with gradient-based fine-tuning
Datasets Pre PMF SoftMerge SMAT Pre+full PMF+full SoftMerge+full SMAT+full Pre+LoRA PMF+LoRA SMAT+LoRA
ImageNet 73.48 73.54 74.33 74.69 73.54 74.59 74.71 75.24 74.22 73.54 75.72
Aircraft 62.17 88.33 88.80 89.78 75.4 88.33 90.60 90.01 80.8 89.75 90.71
Omniglot 54.33 91.79 91.24 89.84 78.7 91.79 92.01 90.83 80.8 92.78 90.99
CUB 85.37 91.02 91.54 92.57 85.4 91.02 91.95 92.57 85.8 91.17 92.57
DTD 83.67 81.64 80.98 86.29 86.9 86.61 86.84 88.41 86.8 86.73 88.28
Quickdraw 60.59 79.23 78.98 79.17 73.6 79.23 79.90 79.17 72.7 79.23 78.83
Fungi 56.26 74.2 72.40 73.31 54.7 74.20 72.40 73.31 59.8 75.44 73.31
VGGFlower 94.45 94.12 96.89 97.22 94.2 94.12 97.01 97.22 94.8 96.05 97.25
ID Avg 71.29 84.23 84.40 85.36 77.81 84.99 85.56 85.84 79.47 85.59 85.88
TrafficSig 53.7 54.37 56.21 57.72 87.3 88.85 89.91 91.33 88.1 89.14 90.18
MSCOCO 54.58 57.04 55.75 58.81 61.5 62.59 62.15 63.11 62.1 61.71 63.38
Cifar10 85.64 80.82 84.58 87.05 92.48 89.61 91.84 92.21 93.33 91.53 92.46
Cifar100 76.86 69.11 70.85 77.46 86.13 82.54 85.88 86.12 86.17 85.06 85.88
MNIST 78.57 93.33 94.16 94.43 92.54 96.44 96.20 96.73 94.98 96.41 96.46
Sketch 47.25 41.10 43.30 47.76 56.39 49.65 53.85 56.67 57.34 47.59 55.63
Pet 91.73 91.37 89.84 91.97 92.03 91.73 90.48 91.97 92.06 92.01 92.31
Clipart 55.19 53.92 54.83 58.97 67.18 62.83 65.50 65.79 66.51 60.6 66.07
Food 62.64 61.89 63.04 65.59 65.08 62.97 63.36 66.99 65.06 62.71 67.77
Cars 34.58 38.00 36.21 36.79 40.98 40.07 41.62 42.39 39.49 42.37 40.05
OOD Avg 64.07 64.10 64.87 67.65 74.16 72.73 74.08 75.32 74.51 72.91 75.02
@@
Evaluation.  At meta-testing time, we resort to the ProtoNet (Snell et al., 2017) classifier for performing direct inference on each few-shot testing tasks without further adaptation. When considering task-specific fine-tuning on the support sets, we follow the same protocols in PMF (Hu et al., 2022) for all models, using Full (fine-tuning the entire model) and LoRA (Hu et al., 2021) as fine-tuning methods. Namely, for each dataset, we perform a hyperparameter search on a few validation tasks to obtain the optimal learning rate for task-specific fine-tuning using the Adam optimizer for 50 steps for each testing tasks from that dataset.
5.1 SMAT achieves new SOTA performance